# Towards Inclusive Communication: A Unified LLM-Based Framework for Sign Language, Lip Movements, and Audio Understanding

### v1

@@ Abstract
Audio is the primary modality for human communication and has driven the success of Automatic Speech Recognition (ASR) technologies. However, such systems remain inherently inaccessible to individuals who are deaf or hard of hearing. Visual alternatives such as sign language and lip reading offer effective substitutes, and recent advances in Sign Language Translation (SLT) and Visual Speech Recognition (VSR) have improved audio-less communication. Yet, these modalities have largely been studied in isolation, and their integration within a unified framework remains underexplored. In this paper, we introduce the first unified framework capable of handling diverse combinations of sign language, lip movements, and audio for spoken-language text generation. We focus on three main objectives: (i) designing a unified, modality-agnostic architecture capable of effectively processing heterogeneous inputs; (ii) exploring the underexamined synergy among modalities, particularly the role of lip movements as non-manual cues in sign language comprehension; and (iii) achieving performance on par with or superior to state-of-the-art models specialized for individual tasks. Building on this framework, we achieve performance on par with or better than task-specific state-of-the-art models across SLT, VSR, ASR, and AVSR. Furthermore, our analysis reveals that explicitly modeling lip movements as a separate modality significantly improves SLT performance. The code and models are available https://bit.ly/UniSLA.
1 Introduction
@@
Audio plays a central role in human communication and serves as the primary modality for conveying linguistic meaning. Accordingly, Automatic Speech Recognition (ASR) technologies have seen rapid advancement  (Amodei et al. 2016; Kim, Hori, and Watanabe 2017; Prabhavalkar et al. 2023), enabling seamless voice-based interaction with smart devices in daily life.
@@
However, audio-based interfaces remain fundamentally inaccessible to individuals who are deaf or hard of hearing. For such users, sign language and lip movements serve as effective visual alternatives. Sign language is a gesture-based linguistic system with its own grammar and syntax, while lip reading allows the interpretation of spoken language through the observation of mouth movements. Recent research has leveraged these alternatives through Sign Language Translation (SLT) (Chen et al. 2024; Jiao, Min, and Chen 2024; Rust et al. 2024; Jang et al. 2025a) and Visual Speech Recognition (VSR) (Ma, Petridis, and Pantic 2022; Kim et al. 2023), making audio-less communication more accessible.
@@
Building on this progress in each area, recent efforts have aimed at integrating multiple modalities into a unified architecture (Makino et al. 2019; Hsu and Shi 2022; Haliassos et al. 2024a). These systems simplify deployment and improve performance by capturing complementary information, as demonstrated in AVSR (Afouras et al. 2018). However, prior work has largely focused on fusing audio and lip movements, overlooking the broader potential of combining sign language, lip movements, and audio. A unified framework that supports flexible combinations of these modalities not only enables more inclusive interaction for users with varying hearing capabilities but also overcomes the limitations of treating each modality in isolation, such as redundant training pipelines, siloed representations, and restricted cross-modal learning. Furthermore, this integration paves the way for more efficient training and richer multimodal representations that reflect the natural synergy between modalities in human communication.
@@
In this paper, we present a unified framework that supports various input combinations of sign language, lip movements, and audio, enabling a single model to generate spoken-language text across a range of communication scenarios, particularly SLT, VSR, ASR, and AVSR. As the first study to jointly investigate these modalities within a single system, we pursue three primary objectives: (i) to design a simple and unified architecture for processing the three modalities, (ii) to explore the previously underexamined synergy between these modalities, particularly by investigating the role of lip movements as non-manual linguistic cues in sign language, and (iii) to achieve performance comparable to task-specific models using a single model. To achieve these goals, we design a system that aligns and integrates heterogeneous modalities into a unified representation. Specifically, we start from the insight that input modalities differ in form but share the same objective of conveying linguistic content. Prior work in AVSR (Zhu et al. 2023; Yeo et al. 2025) has shown that although lip movements and audio differ in granularity, such as phonemes versus visemes (i.e., the visual counterparts of phonemes), their frame-level features can be temporally aligned and jointly leveraged through simple fusion strategies. Building on this observation, we hypothesize that frame-level sign features, which contain gloss information (i.e., representations of individual sign meanings), can likewise be temporally aligned and integrated. Based on this hypothesis, we model a unified linguistic representation by aligning and fusing the temporally synchronized features from the three modalities. These unified representations are then passed to a Large Language Model (LLM) decoder, leveraging recent findings that LLMs are capable of context-aware decoding and cross-modal linguistic interpretation (Gong et al. 2024; Cappellazzo et al. 2025).
@@
Beyond its practical utility, this unified setting enables analytical insights into the contribution of individual modalities. In particular, it allows us to systematically examine the impact of incorporating lip movements into SLT as a distinct modality, rather than treating them as part of general visual input. Although non-manual cues in sign language, especially lip movements, are known to convey grammatical and semantic distinctions (Sandler 2009), they have been underutilized in prior work due to the absence of architectural frameworks that treat them separately. Our unified model reveals that leveraging lip movements as a complementary modality significantly improves translation quality, highlighting their essential role in sign language communication.
@@
  • Unified Tri-Modal Architecture We propose a single, unified framework that flexibly integrates sign language, lip movements, and audio, supporting various modality combinations for spoken-language generation tasks (SLT, VSR, ASR, AVSR) within a single model
  • Linguistic Role of Lip Movements in SLT We demonstrate that explicitly modeling lip movements as a separate modality significantly enhances sign language translation, highlighting their value as essential non-manual cues in sign language.
  • Two-stage multi-task learning strategy. We introduce a two-stage learning schedule that first focuses on visual tasks, then jointly optimizes all objectives, improving SLT and VSR while preserving ASR and AVSR performance.
2 Related Works
2.1 Automatic Speech Recognition
@@
Advances in deep learning architectures have played a crucial role in improving ASR performance. Early studies replaced traditional HMM-GMM acoustic models with Deep Neural Networks (DNNs) (Mohamed, Dahl, and Hinton 2011; Hinton et al. 2012). To further capture local time-frequency patterns, Convolutional Neural Networks (CNNs) (Abdel-Hamid et al. 2012, 2014) were introduced, while Recurrent Neural Networks (RNNs) (Graves, Mohamed, and Hinton 2013) model sequential dependencies. More recently, attention-based Transformer (Vaswani 2017) architectures have enabled the integration of long-range context, contributing to substantial performance gains.
@@
In addition to architectural advances, the availability of large-scale datasets (Panayotov et al. 2015; Ardila et al. 2019) and improved learning methods have also played key roles in the progress of ASR. Techniques such as Connectionist Temporal Classification (CTC) (Graves et al. 2006) and sequence-to-sequence (Sutskever 2014) learning have enabled end-to-end speech recognition. While supervised approaches require large labeled datasets, self-supervised learning methods (Baevski et al. 2020; Hsu et al. 2021) have been developed to leverage vast amounts of unlabeled audio. In parallel, Whisper (Radford et al. 2023), trained on 680,000 hours of multilingual and noisy speech via large-scale supervised learning, has emerged as a powerful foundation model for speech recognition. With the rise of powerful audio encoders, researchers (Rubenstein et al. 2023; Chu et al. 2023) have begun integrating LLMs (Achiam et al. 2023; Touvron et al. 2023) into ASR pipelines, enhancing contextual understanding and recognition quality.
2.2 Visual Speech Recognition
@@
Despite the importance of lip movements for individuals who are deaf or hard of hearing, Visual Speech Recognition (VSR) received comparatively less attention than ASR in its early stages, primarily due to the difficulty of collecting and annotating large-scale video data. Recent releases of large-scale datasets (Afouras, Chung, and Zisserman 2018; Chung and Zisserman 2016) have accelerated significant progress in VSR research.
@@
VSR architectures have followed a development path broadly similar to ASR, evolving from CNNs (Noda et al. 2014) for spatiotemporal feature extraction, to RNNs (Petridis, Li, and Pantic 2017) for temporal modeling, and more recently to Transformer-based models (Afouras et al. 2018) that capture long-range dependencies. However, VSR poses unique challenges. A notable example is the prevalence of homophenes (i.e., different phonemes that appear visually similar) (Kim, Yeo, and Ro 2022), which introduce ambiguity and limit performance. To overcome these challenges, multimodal approaches that incorporate audio signals as auxiliary supervision have become increasingly popular, leveraging the temporal alignment and rich acoustic cues of audio. These include knowledge distillation from pretrained ASR teachers (Ren et al. 2021), and self-supervised learning using audio-visual speech units as pseudo labels (Shi et al. 2022; Haliassos et al. 2022), and pseudo-labeling approaches (Ma et al. 2023; Yeo et al. 2024b) that generate text transcriptions using pretrained ASR models. Building on these advances, recent work (Yeo et al. 2024a; Cappellazzo et al. 2025) has explored integrating LLMs into VSR pipelines, enabling contextual reasoning over visually ambiguous inputs and improving robustness in real-world scenarios.
2.3 Sign Language Translation
@@
Sign Language Translation (SLT) builds on prior tasks like Isolated and Continuous Sign Language Recognition (ISLR, CSLR), which extract gloss representations from sign videos. These areas have advanced significantly thanks to datasets for ISLR (Li et al. 2020; Joze and Koller 2018) and CSLR (Koller, Forster, and Ney 2015; Zhou et al. 2021), enabling large-scale deep learning models (Ahn, Jang, and Chung 2024; Jang et al. 2023). Both tasks use deep spatiotemporal architectures like I3D (Carreira and Zisserman 2017) and Video Swin Transformers (Liu et al. 2022), but differ in objectives: ISLR classifies segmented clips via cross-entropy, while CSLR (Chen et al. 2022; Jang et al. 2022) handles unsegmented input using gloss-level supervision and CTC loss (Graves et al. 2006).
@@
With the availability of SLT datasets (Duarte et al. 2021) and visual representations learned through ISLR and CSLR, SLT research has seen significant progress. While early approaches primarily relied on visual backbones pretrained on CSLR tasks, recent work has increasingly focused on gloss-free SLT methods. This shift is largely driven by the growing difficulty of obtaining manual gloss annotations as datasets scale (Uthus, Tanzer, and Georg 2023). In response, two main gloss-free approaches have emerged to effectively train the visual backbone without relying on gloss annotations: contrastive learning for visual-text alignment (Jiao, Min, and Chen 2024; Lin et al. 2023; Zhou et al. 2023; Ye et al. 2024), and pseudo-gloss generation for weak supervision (Prajwal et al. 2022; Guo, Li, and Cohn 2025). In parallel, increasing attention has been directed toward improving the decoding stage through the integration of LLMs (Gong et al. 2024; Wong, Camgoz, and Bowden 2024). Since gloss sequences differ grammatically from natural language, LLM-based decoders offer clear advantages in generating fluent and natural spoken-language text. Building on the strengths of LLM-based decoders, more recent work (Jang et al. 2025b) leverages contextual cues such as preceding sentences or background descriptions.
@@
2.4 Unified Modeling across Modalities
Unified models for performing multi-task VSR, ASR, and AVSR have been explored in several studies. One early approach was based on RNN-T (Makino et al. 2019), trained on a large-scale audio-visual dataset (31k hours) and incorporating strategies such as modality dropout during training. However, its performance significantly lagged behind that of task-specific models. AV-HuBERT (Shi et al. 2022) was the first to achieve performance comparable to task-specific models by leveraging a single encoder trained with self-supervised learning and modality dropout. However, it still relied on task-specific decoders. Building on this encoder, u-HuBERT (Hsu and Shi 2022) extended the modality dropout strategy to the fine-tuning stage, enabling the use of a single encoder-decoder architecture and achieving strong performance across all tasks. More recently, USR (Haliassos et al. 2024a) introduced a semi-supervised learning strategy that utilizes online pseudo-labels generated by an exponential moving average (EMA)-based teacher during the fine-tuning stage. This approach demonstrated its effectiveness by achieving state-of-the-art performance on LRS3.
@@ 3.2 Modality Encoders
Sign Encoder
To generate sign features that capture gloss information, we adopt a sign encoder based on the Video Swin Transformer, following (Jang et al. 2025b). The encoder is pre-trained on the ISLR task, which aims to predict gloss annotations from short signing video clips. Given this pre-trained sign encoder and input sign videos $X_{s}=\{x_{1},...,x_{T_{s}}\}$, where $T_{s}$ denotes the number of video frames, the sign encoder $E_{s}$ extracts a sequence of sign features by capturing spatiotemporal patterns from the signing video. The resulting feature sequence is denoted as $F_{s}=E_{s}(X_{s})$, where $F_{s}\in\mathbb{R}^{T_{s}\times D_{s}}$ represents sign features with the same temporal resolution as the input.
Lip Encoder
Following (Yeo et al. 2024a), which demonstrates that lip features generated by self-supervised models can be effectively aligned with LLMs, we employ AV-HuBERT (Shi et al. 2022) as our lip encoder. Given an input lip video $X_{v}=\{x_{1},...,x_{T_{v}}\}$, where $T_{v}$ denotes the number of frames, the lip encoder $E_{v}$ extracts a sequence of lip features that capture viseme information by modeling the temporal dynamics of lip movements. The resulting feature sequence is denoted as $F_{v}=E_{v}(X_{v})$, where $F_{v}\in\mathbb{R}^{T_{v}\times D_{v}}$ represents lip features with the same temporal resolution as the input.
Audio Encoder
Similarly, motivated by recent studies (Chu et al. 2023) showing that audio features extracted by Whisper (Radford et al. 2023) can also be effectively aligned with LLMs, we adopt it as our audio encoder. Given a log-Mel spectrogram input $X_{a}=\{x_{1},...,x_{T_{a}^{\prime}}\}$, where $T_{a}^{\prime}$ denotes the number of audio frames and each $x_{t}\in\mathbb{R}^{B}$ is a feature vector of $B$ Mel-frequency bins, the audio encoder $E_{a}$ extracts a sequence of audio features that encode phoneme information. The resulting feature sequence is denoted as $F_{a}=E_{a}(X_{a})$, where $F_{a}\in\mathbb{R}^{T_{a}\times D_{a}}$ represents audio features. Note that $T_{a}^{\prime}$ differs from $T_{a}$ due to internal downsampling applied within the Whisper encoder.
3.3 Modality Alignment & Fusion
Length Adapters

### v2

# Towards Inclusive Communication: A Unified Framework for Generating Spoken Language from Sign, Lip, and Audio

@@ Abstract
Audio is the primary modality for human communication and has driven the success of Automatic Speech Recognition (ASR) technologies. However, such audio-centric systems inherently exclude individuals who are deaf or hard of hearing. Visual alternatives such as sign language and lip reading offer effective substitutes, and recent advances in Sign Language Translation (SLT) and Visual Speech Recognition (VSR) have improved audio-less communication. Yet, these modalities have largely been studied in isolation, and their integration within a unified framework remains underexplored. In this paper, we propose the first unified framework capable of handling diverse combinations of sign language, lip movements, and audio for spoken-language text generation. We focus on three main objectives: (i) designing a unified, modality-agnostic architecture capable of effectively processing heterogeneous inputs; (ii) exploring the underexamined synergy among modalities, particularly the role of lip movements as non-manual cues in sign language comprehension; and (iii) achieving performance on par with or superior to state-of-the-art models specialized for individual tasks. Building on this framework, we achieve performance on par with or better than task-specific state-of-the-art models across SLT, VSR, ASR, and Audio-Visual Speech Recognition. Furthermore, our analysis reveals a key linguistic insight: explicitly modeling lip movements as a distinct modality significantly improves SLT performance by capturing critical non-manual cues.
I Introduction
@@
Audio plays a central role in human communication and serves as the primary modality for conveying linguistic meaning. Accordingly, Automatic Speech Recognition (ASR) technologies have seen rapid advancement  [1, 2, 3], enabling seamless voice-based interaction with smart devices in daily life.
@@
However, audio-based interfaces remain fundamentally inaccessible to individuals who are deaf or hard of hearing. For such users, sign language and lip movements serve as effective visual alternatives. Sign language is a gesture-based linguistic system with its own grammar and syntax, while lip reading allows the interpretation of spoken language through the observation of mouth movements. Recent research has leveraged these alternatives through Sign Language Translation (SLT) [4, 5, 6, 7, 8] and Visual Speech Recognition (VSR) [9, 10, 11], making audio-less communication more accessible.
@@
Building on this progress in each area, recent efforts have aimed at integrating multiple modalities into a unified architecture [12, 13, 14]. While these efforts mark a step towards unification, they have been predominantly limited to the audio-lip fusion seen in Audio-Visual Speech Recognition (AVSR) [15, 16]. This narrow focus overlooks the broader potential of combining sign language, lip movements, and audio. A unified framework that supports flexible combinations of these modalities not only enables more inclusive interaction for users with varying hearing capabilities but also overcomes the limitations of treating each modality in isolation, such as redundant training pipelines, siloed representations, and restricted cross-modal learning. Furthermore, this integration paves the way for more efficient training and richer multimodal representations that reflect the natural synergy between modalities in human communication.
@@
In this paper, we present a unified framework that supports various input combinations of sign language, lip movements, and audio, enabling a single model to generate spoken-language text across a range of communication scenarios, particularly SLT, VSR, ASR, and AVSR. As the pioneering study to jointly investigate these modalities within a single system, we pursue three primary objectives: (i) to design a simple and unified architecture for processing the three modalities, (ii) to explore the previously underexamined synergy between these modalities, particularly by investigating the role of lip movements as non-manual linguistic cues in sign language, and (iii) to achieve performance comparable to task-specific models using a single model. To achieve these goals, we design a system that aligns and integrates heterogeneous modalities into a unified representation. Specifically, we start from the insight that input modalities differ in form but share the same objective of conveying linguistic content. Prior work in AVSR [17, 18] has shown that although lip movements and audio differ in granularity, such as phonemes versus visemes (i.e., the visual counterparts of phonemes), their frame-level features can be temporally aligned and jointly leveraged through simple fusion strategies. Building on this observation, we hypothesize that frame-level sign features, which contain gloss information (i.e., representations of individual sign meanings), can likewise be temporally aligned and integrated. Based on this hypothesis, we model a unified linguistic representation by aligning and fusing the temporally synchronized features from the three modalities. These unified representations are then passed to a Large Language Model (LLM) decoder, leveraging recent findings that LLMs are capable of context-aware decoding and cross-modal linguistic interpretation [19, 20].
@@
Beyond its practical utility, this unified setting enables analytical insights into the contribution of individual modalities. In particular, it allows us to systematically examine the impact of incorporating lip movements into SLT as a distinct modality, rather than treating them as part of general visual input. Although non-manual cues in sign language, especially lip movements, are known to convey grammatical and semantic distinctions [21], they have been computationally underutilized in prior work, primarily due to the lack of architectural frameworks capable of treating them as a distinct input stream. Our work directly addresses this architectural gap. The proposed model reveals that leveraging lip movements as a complementary modality significantly improves translation quality, highlighting their essential role in sign language communication.
@@
  • A Unified Tri-Modal Framework A single architecture that flexibly integrates sign language, lip movements, and audio to perform SLT, VSR, ASR, and AVSR within one model.
  • Novel Linguistic Insight on SLT The first empirical evidence that explicitly modeling lip movements as a separate modality significantly enhances sign language translation by capturing non-manual cues.
  • An Effective Two-Stage Training Strategy A specialized training schedule that resolves the learning imbalance between visual and audio tasks, optimizing performance across all modalities.
II Related Works
II-A Automatic Speech Recognition
@@
Advances in deep learning architectures have played a crucial role in improving ASR performance. Early studies replaced traditional HMM-GMM acoustic models with Deep Neural Networks (DNNs) [22]. To further capture local time-frequency patterns, Convolutional Neural Networks (CNNs) [23] were introduced, while Recurrent Neural Networks (RNNs) [24] model sequential dependencies. More recently, attention-based Transformer [25] architectures have enabled the integration of long-range context, contributing to substantial performance gains.
@@
In addition to architectural advances, the availability of large-scale datasets [26] and improved learning methods have also played key roles in the progress of ASR. Techniques such as Connectionist Temporal Classification (CTC) [27] and sequence-to-sequence [28] learning have enabled end-to-end speech recognition. While supervised approaches require large labeled datasets, self-supervised learning methods [29, 30] have been developed to leverage vast amounts of unlabeled audio. In parallel, Whisper [31], trained on 680,000 hours of multilingual and noisy speech via large-scale supervised learning, has emerged as a powerful foundation model for speech recognition. With the rise of powerful audio encoders, researchers [32] have begun integrating LLMs [33, 34] into ASR pipelines, enhancing contextual understanding and recognition quality.
II-B Visual Speech Recognition
@@
Despite the importance of lip movements for individuals who are deaf or hard of hearing, Visual Speech Recognition (VSR) received comparatively less attention than ASR in its early stages, primarily due to the difficulty of collecting and annotating large-scale video data. Recent releases of large-scale datasets [35, 36, 37, 38] have accelerated significant progress in VSR research.
@@
VSR architectures have followed a development path broadly similar to ASR, evolving from CNNs [39] for spatiotemporal feature extraction, to RNNs [40] for temporal modeling, and more recently to Transformer-based models [15] that capture long-range dependencies. However, VSR poses unique challenges. A notable example is the prevalence of homophenes (i.e., different phonemes that appear visually similar) [41], which introduce ambiguity and limit performance. To overcome these challenges, multimodal approaches that incorporate audio signals as auxiliary supervision have become increasingly popular, leveraging the temporal alignment and rich acoustic cues of audio. These include knowledge distillation from pretrained ASR teachers [42], and self-supervised learning using audio-visual speech units as pseudo labels [43, 44], and pseudo-labeling approaches [45, 46] that generate text transcriptions using pretrained ASR models. Building on these advances, recent work [47, 20] has explored integrating LLMs into VSR pipelines, enabling contextual reasoning over visually ambiguous inputs and improving robustness in real-world scenarios.
II-C Sign Language Translation
@@
Sign Language Translation (SLT) builds on prior tasks like Isolated and Continuous Sign Language Recognition (ISLR, CSLR), which extract gloss representations from sign videos. These areas have advanced significantly thanks to datasets for ISLR [48] and CSLR [49], enabling large-scale deep learning models [50, 51, 52, 53]. Both tasks use deep spatiotemporal architectures like I3D [54] and Video Swin Transformers [55], but differ in objectives: ISLR classifies segmented clips via cross-entropy, while CSLR [56, 57] handles unsegmented input using gloss-level supervision and CTC loss [27].
@@
With the availability of SLT datasets [58] and visual representations learned through ISLR and CSLR, SLT research has seen significant progress. While early approaches primarily relied on visual backbones pretrained on CSLR tasks, recent work has increasingly focused on gloss-free SLT methods. This shift is largely driven by the growing difficulty of obtaining manual gloss annotations as datasets scale [59]. In response, two main gloss-free approaches have emerged to effectively train the visual backbone without relying on gloss annotations: contrastive learning for visual-text alignment [5, 60, 61], and pseudo-gloss generation for weak supervision [62, 63]. In parallel, increasing attention has been directed toward improving the decoding stage through the integration of LLMs [19, 64]. Since gloss sequences differ grammatically from natural language, LLM-based decoders offer clear advantages in generating fluent and natural spoken-language text. Building on the strengths of LLM-based decoders, more recent work [65] leverages contextual cues such as preceding sentences or background descriptions.
@@
II-D Unified Modeling across Modalities
Unified models for performing multi-task VSR, ASR, and AVSR have been explored in several studies. One early approach was based on RNN-T [12], trained on a large-scale audio-visual dataset (31k hours) and incorporating strategies such as modality dropout during training. However, its performance significantly lagged behind that of task-specific models. AV-HuBERT [43] was the first to achieve performance comparable to task-specific models by leveraging a single encoder trained with self-supervised learning and modality dropout. However, it still relied on task-specific decoders. Building on this encoder, u-HuBERT [13] extended the modality dropout strategy to the fine-tuning stage, enabling the use of a single encoder-decoder architecture and achieving strong performance across all tasks. More recently, USR [14] introduced a semi-supervised learning strategy that utilizes online pseudo-labels generated by an exponential moving average (EMA)-based teacher during the fine-tuning stage. This approach demonstrated its effectiveness by achieving state-of-the-art performance on LRS3.
@@ III-B Modality Encoders
III-B1 Sign Encoder
To generate sign features that capture gloss information, we adopt a sign encoder based on the Video Swin Transformer, following [65]. The encoder is pre-trained on the ISLR task, which aims to predict gloss annotations from short signing video clips. Given this pre-trained sign encoder and input sign videos $X_{s}=\{x_{1},...,x_{T_{s}}\}$, where $T_{s}$ denotes the number of video frames, the sign encoder $E_{s}$ extracts a sequence of sign features by capturing spatiotemporal patterns from the signing video. The resulting feature sequence is denoted as $F_{s}=E_{s}(X_{s})$, where $F_{s}\in\mathbb{R}^{T_{s}\times D_{s}}$ represents sign features with the same temporal resolution as the input.
III-B2 Lip Encoder
Following [47], which demonstrates that lip features generated by self-supervised models can be effectively aligned with LLMs, we employ AV-HuBERT [43] as our lip encoder. Given an input lip video $X_{v}=\{x_{1},...,x_{T_{v}}\}$, where $T_{v}$ denotes the number of frames, the lip encoder $E_{v}$ extracts a sequence of lip features that capture viseme information by modeling the temporal dynamics of lip movements. The resulting feature sequence is denoted as $F_{v}=E_{v}(X_{v})$, where $F_{v}\in\mathbb{R}^{T_{v}\times D_{v}}$ represents lip features with the same temporal resolution as the input.
III-B3 Audio Encoder
Similarly, motivated by recent studies [32] showing that audio features extracted by Whisper [31] can also be effectively aligned with LLMs, we adopt it as our audio encoder. Given a log-Mel spectrogram input $X_{a}=\{x_{1},...,x_{T_{a}^{\prime}}\}$, where $T_{a}^{\prime}$ denotes the number of audio frames and each $x_{t}\in\mathbb{R}^{B}$ is a feature vector of $B$ Mel-frequency bins, the audio encoder $E_{a}$ extracts a sequence of audio features that encode phoneme information. The resulting feature sequence is denoted as $F_{a}=E_{a}(X_{a})$, where $F_{a}\in\mathbb{R}^{T_{a}\times D_{a}}$ represents audio features. Note that $T_{a}^{\prime}$ differs from $T_{a}$ due to internal downsampling applied within the Whisper encoder.
III-C Modality Alignment & Fusion
III-C1 Length Adapters